writing an equation, in standard form, of the ellipse with the center at (-1,3), vertex at (3,3), and a minor axis of length 2.
I have this so far (x+1)^2/( )+ (y-3)^2/( )=1
I don't know what to do with the minor axis???

Answer by lwsshak3(11628)

writing an equation, in standard form, of the ellipse with the center at (-1,3), vertex at (3,3), and a minor axis of length 2.
I have this so far (x+1)^2/( )+ (y-3)^2/( )=1
I don't know what to do with the minor axis?
**
Standard form of an equation for an ellipse with horizontal major axis:
(x-h)^2/a^2+(y-k)^2/b^2=1, (a>b), (h,k) being the (x,y) coordinates of the center.
For given equation of ellipse:
center:(-1, 3) (given)
..
distance from center to vertex on major axis, -1 to 3=4=a
a^2=16
..
length of minor axis=2b=2
b=1
b^2=1
Equation:
(x+1)^2/16+(y-3)^2/1=1
..
When doing conics like ellipses, parabolas and hyperbolas, you won't get stuck if you can remember the standard form for the equation of each conic. For ellipses, remember one form for ellipses with horizontal major axis and one other form with vertical major axis.